# Bose-Einstein condensation of THz photons in an optical microcavity with Landau-quantized electrons

### v1

@@ A.2 Electron-photon
$\hat{H}_{e-ph}{=}\frac{e}{m^{*}c}\!\!\int\!\!d^{3}\bm{r}\hat{\psi}^{\dagger}(\bm{r})\!\!\left[\!\hat{\bm{A}}(\bm{r})\!\left(\hat{\bm{p}}{-}\frac{eH}{c}y\bm{e}_{x}\!\right)\!\right]\!\hat{\psi}_{(}\bm{r}).$
@@
×${\times}\left[\frac{\epsilon_{\bm{q},\xi}^{x}{-}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{+}1}{+}\frac{\epsilon_{\bm{q},\xi}^{x}{+}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{-}1}\right]$
@@
×${\times}\left[\frac{\epsilon_{\bm{q},\xi}^{x*}{-}i\epsilon^{y*}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{+}1}{+}\frac{\epsilon_{\bm{q},\xi}^{x*}{+}i\epsilon^{y*}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{-}1}\right]$
@@
×${\times}\!\left[\!\!\left(\frac{\epsilon_{\bm{q},\xi}^{x}{\pm}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\right)^{\pm}\!\!\!\!\!\delta_{m,n{+}1}{+}\!\left(\frac{\epsilon_{\bm{q},\xi}^{x}{\mp}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\right)^{\pm}\!\!\!\!\!\delta_{m,n{-}1}\right].$ (30)
@@
×$\times\left(\frac{|\bm{q}(l)|\mp\xi q_{z}(l)}{2|\bm{q}(l)|}\delta_{m,n+1}+\frac{|\bm{q}(l)|\pm\xi q_{z}(l)}{2|\bm{q}(l)|}\delta_{m,n-1}\right),$
@@ Appendix B The impact of magnetic field gradient
[…]
@@ Appendix C Calculation details for condensate mode kinetics
Here we address all the kinetic processes to derive expressions for the rates $\nu_{\rm pump}$ (7), $\nu_{\rm reverse}$ (8), $\nu_{\rm loss}$ (9) and $\nu_{\rm satell}$ (12) introduced in the main text.
@@
$\nu_{\rm pump}$ ${=}\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{-}\frac{\omega_{\rm cut}}{\omega_{\rm pump}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{q}^{2}}\!\right)$
×$\times\sum_{n,s}(n+1)(n+2)\left(N^{\rm el}_{n,s}-N^{\rm el}_{n+2,s}\right),$ (41)
$\nu_{\rm reverse}$ ${=}\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{-}\frac{\omega_{\rm cut}}{\omega_{\rm pump}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{q}^{2}}\!\right)$
×$\times\sum_{n,s}(n+1)(n+2)N^{\rm el}_{n+2,s}\left(1-N^{\rm el}_{n,s}\right).$ (42)
@@
$\nu_{\rm loss}{=}\frac{SN_{\rm layer}}{2\pi l_{H}^{2}}\!\sum_{n}\!\frac{2\pi}{\hbar}\!\!\int\!\frac{d^{3}kV}{(2\pi\hbar)^{3}}|A_{n,m}^{\rm loss}|^{2}\!\!\left(\!\frac{\delta(\hbar\omega_{c}{-}\hbar\omega_{s}{-}(m{-}n)\hbar\omega_{H}{-}ku)}{1{-}\exp(-ku/T)}{+}\frac{\delta(\hbar\omega_{c}{-}\hbar\omega_{s}{-}(m{-}n)\hbar\omega_{H}{+}ku)}{\exp(ku/T){-}1}\right)\left(N^{\rm el}(E_{n}^{+}){-}N^{\rm el}(E_{m}^{-})\right).$
@@
where $k<0$/$k>0$ stands for phonon absorption/emission, $\Delta_{m}=\hbar\omega_{c}-\hbar\omega_{s}-m\hbar\omega_{H}$ is the photon energy offset from resonant absorption frequency.
@@
$\nu_{\rm loss}(\hbar\omega_{c})$ $=\omega_{H}\frac{D_{e}^{2}\hbar\omega_{H}T}{2\pi\rho u^{2}(\hbar u)^{3}}\frac{e^{2}\psi_{\rm cond}^{2}N_{\rm layer}}{\varepsilon\hbar\omega_{c}}\sum_{n,m}\left(\frac{\Delta_{m}}{\hbar\omega_{r}+\hbar\omega_{c}}\right)^{2}\left(N^{\rm el}(E_{n}^{+}){-}N^{\rm el}(E_{m}^{-})\right)\times$ (44)
×$\left.\times\int_{0}^{\pi}\frac{d\cos\theta}{4}\left(\frac{\Delta_{m}l_{H}\sin\theta}{\hbar u}\right)^{2}V_{00}^{2}\left(\frac{\Delta_{m}^{2}l_{z}^{2}}{2\hbar^{2}u^{2}}\cos^{2}\theta\right)V_{nm}^{2}\left(\frac{\Delta_{m}^{2}l_{H}^{2}}{2\hbar^{2}u^{2}}\sin^{2}\theta\right)\right|_{\Delta_{m}=\hbar\omega_{c}-\hbar\omega_{s}-(m-n)\hbar\omega_{H}}$ (45)
@@
In the same fashion as for the $\nu_{pump}$ and $\nu_{\rm reverse}$, one may consider the interaction with the satellite modes:
$\nu_{\rm satell}$ $=\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}{\rm electron}\\ {\rm photon}\end{subarray}}\!\!\!\left(\frac{\lambda_{n-s,n}^{-}(q_{\rm satell},l)\lambda^{c+}_{n,n-s}}{E^{s}_{n}+\hbar\omega_{c}-E_{n-s}^{-s}}\right)^{2}$
×$\times\delta(\hbar\omega_{\rm satell}-\hbar\omega_{c})N^{\rm el}_{T}(n^{s})\left(1-N^{\rm el}_{T}(n^{s})\right).$ (46)
@@
$\nu_{\rm satell}$ $=\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{+}\frac{\omega_{\rm cut}}{\omega_{c}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{c}^{2}}\right)$
×$\times\sum_{n,s}\left({\rm max}(n,n-s)\right)^{2}N^{\rm el}_{T}(n^{s})\left(1-N^{\rm el}_{T}(n^{s})\right).$ (47)
@@ Appendix E Conditions for neglecting the nonlinearity
$\Sigma_{n+1}^{+}$ $=\frac{\hbar\omega_{r}}{2}\left(\sqrt{1+\left|\frac{2\lambda_{n+1,n}(\bm{0})}{\hbar\omega_{r}}\right|^{2}N_{0}}-1\right),$ (52)
$\Sigma_{n}^{-}$ $=-\frac{\hbar\omega_{r}}{2}\left(\sqrt{1+\left|\frac{2\lambda_{n+1,n}(\bm{0})}{\hbar\omega_{r}}\right|^{2}N_{0}}-1\right).$ (53)
@@
$n_{0}^{\rm crit}(n)\sim\frac{\varepsilon\hbar u}{2\pi ne^{2}\psi_{c}^{2}l_{H}^{3}}\sim\frac{1}{n}(10^{13}-10^{14}){\rm cm}^{-2}.$ (56)
@@
$W{\simeq}\frac{c}{\sqrt{\varepsilon}}\int_{\hbar\omega_{\rm pump}^{\rm min}(q)}^{\hbar\omega_{\rm pump}^{\rm max}(q)}\!\frac{d^{2}q}{(2\pi\hbar)^{3}}N_{\rm pump}\left(\frac{E_{q}}{T_{\rm pump}}\right)\simeq$
$\simeq\frac{\varepsilon\omega_{\rm pump}\Delta_{\rm pump}T_{\rm pump}}{2\pi^{2}(\hbar c)^{2}}<Q\nu\frac{\delta H}{H}\frac{\hbar\omega_{H}T_{\rm pump}}{(\hbar c)^{2}}$ (57)
@@
$n_{0}^{\rm BB}<\frac{W}{\nu_{0}+\nu}<Q\frac{\delta H}{H}\frac{\hbar\omega_{H}T_{\rm pump}}{(\hbar c)^{2}}\sim 10^{5}Q\ {\rm cm}^{-2}.$ (58)
We can compare the critical density and photon density achieved by blackbody pumping: $n_{0}^{\rm BB}<10^{5}Q{\rm cm}^{-2}<10^{11}{\rm cm}^{-2}<n_{0}^{\rm crit}$. Thus, nonlinearities are safely negligible for our system.

### v2

@@ A.2 Electron-photon
$\hat{H}_{e-ph}{=}\frac{e}{m^{*}_{e}c}\!\!\int\!\!d^{3}\bm{r}\hat{\psi}^{\dagger}(\bm{r})\!\!\left[\!\hat{\bm{A}}(\bm{r})\!\left(\hat{\bm{p}}{-}\frac{eH}{c}y\bm{e}_{x}\!\right)\!\right]\!\hat{\psi}_{(}\bm{r}).$
@@
${\times}\left[\frac{\epsilon_{\bm{q},\xi}^{x}{-}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{+}1}{+}\frac{\epsilon_{\bm{q},\xi}^{x}{+}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{-}1}\right]$
@@
${\times}\left[\frac{\epsilon_{\bm{q},\xi}^{x*}{-}i\epsilon^{y*}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{+}1}{+}\frac{\epsilon_{\bm{q},\xi}^{x*}{+}i\epsilon^{y*}_{\bm{q},\xi}}{\sqrt{2}}\delta_{m,n{-}1}\right]$
@@
${\times}\!\left[\!\!\left(\frac{\epsilon_{\bm{q},\xi}^{x}{\pm}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\right)^{\pm}\!\!\!\!\!\delta_{m,n{+}1}{+}\!\left(\frac{\epsilon_{\bm{q},\xi}^{x}{\mp}i\epsilon^{y}_{\bm{q},\xi}}{\sqrt{2}}\right)^{\pm}\!\!\!\!\!\delta_{m,n{-}1}\right].$ (34)
@@
$\times\left(\frac{|\bm{q}(l)|\mp\xi q_{z}(l)}{2|\bm{q}(l)|}\delta_{m,n+1}+\frac{|\bm{q}(l)|\pm\xi q_{z}(l)}{2|\bm{q}(l)|}\delta_{m,n-1}\right),$
@@ Appendix B The impact of magnetic field gradient
$f(H)$ $=\langle f(H)\rangle\equiv\int_{-L/2}^{L/2}f\left(H(1+\kappa x/L)\right)dx/L.$ (40)
@@ Appendix C Calculation details for condensate mode kinetics
Here we address all the kinetic processes to derive expressions for the rates $R$ (7), $r$ (8), $\varkappa_{\rm loss}$ (9) and $\varkappa_{\rm sat}$ (13) introduced in the main text.
@@
$R$ ${=}\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{-}\frac{\omega_{\rm cut}}{\omega_{\rm res}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{q}^{2}}\!\right)$
$\times\sum_{n,s}(n+1)(n+2)N^{\rm el}_{n,s}\left(1-N^{\rm el}_{n+2,s}\right),$ (45)
$r$ ${=}\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{-}\frac{\omega_{\rm cut}}{\omega_{\rm res}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{q}^{2}}\!\right)$
$\times\sum_{n,s}(n+1)(n+2)N^{\rm el}_{n+2,s}\left(1-N^{\rm el}_{n,s}\right).$ (46)
@@
$\varkappa_{\rm loss}{=}\frac{SN_{\rm layer}}{2\pi l_{H}^{2}}\!\sum_{n}\!\frac{2\pi}{\hbar}\!\!\int\!\frac{d^{3}kV}{(2\pi\hbar)^{3}}|A_{n,m}^{\rm loss}|^{2}\!\!\left(\!\frac{\delta(\hbar\omega_{c}{-}\hbar\omega_{s}{-}(m{-}n)\hbar\omega_{H}{-}ku)}{1{-}\exp(-ku/T)}{+}\frac{\delta(\hbar\omega_{c}{-}\hbar\omega_{s}{-}(m{-}n)\hbar\omega_{H}{+}ku)}{\exp(ku/T){-}1}\right)\left(N^{\rm el}(E_{n}^{+}){-}N^{\rm el}(E_{m}^{-})\right).$
@@
where k<0k<0/$k>0$ stands for phonon absorption/emission, $\Delta_{m}=\hbar\omega_{c}-\hbar\omega_{s}-m\hbar\omega_{H}$ is the photon energy offset from resonant absorption frequency.
@@
$\varkappa_{\rm loss}(\hbar\omega_{c})$ $=\omega_{H}\frac{D_{e}^{2}\hbar\omega_{H}T}{2\pi\rho u^{2}(\hbar u)^{3}}\frac{e^{2}\psi_{\rm cond}^{2}N_{\rm layer}}{\varepsilon\hbar\omega_{c}}\sum_{n,m}\left(\frac{\Delta_{m}}{\hbar\omega_{r}+\hbar\omega_{c}}\right)^{2}\left(N^{\rm el}(E_{n}^{+}){-}N^{\rm el}(E_{m}^{-})\right)\times$ (48)
$\left.\times\int_{0}^{\pi}\frac{d\cos\theta}{4}\left(\frac{\Delta_{m}l_{H}\sin\theta}{\hbar u}\right)^{2}V_{00}^{2}\left(\frac{\Delta_{m}^{2}l_{z}^{2}}{2\hbar^{2}u^{2}}\cos^{2}\theta\right)V_{nm}^{2}\left(\frac{\Delta_{m}^{2}l_{H}^{2}}{2\hbar^{2}u^{2}}\sin^{2}\theta\right)\right|_{\Delta_{m}=\hbar\omega_{c}-\hbar\omega_{s}-(m-n)\hbar\omega_{H}}$ (49)
@@
In the same fashion as for the $R$ and $r$, one may consider the interaction with the satellite modes:
$\varkappa_{\rm sat}$ $=\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}{\rm electron}\\ {\rm photon}\end{subarray}}\!\!\!\left(\frac{\lambda_{n-s,n}^{-}(q_{\rm sat},l)\lambda^{c+}_{n,n-s}}{E^{s}_{n}+\hbar\omega_{c}-E_{n-s}^{-s}}\right)^{2}$
$\times\delta(\hbar\omega_{\rm sat}-\hbar\omega_{c})N^{\rm el}_{T}(n^{s})\left(1-N^{\rm el}_{T}(n^{s})\right).$ (50)
@@
$\varkappa_{\rm sat}$ $=\omega_{\rm cut}\alpha N_{\rm layer}\frac{e^{2}\psi_{\rm cond}^{2}}{\sqrt{\varepsilon}\hbar\omega_{c}}\!\!\left(\frac{l_{H}\omega_{H}^{2}}{c(\omega_{c}{+}\omega_{r})}\right)^{\!\!2}\!\!\!\left(1{+}\frac{\omega_{\rm cut}}{\omega_{c}}{+}\frac{\omega_{\rm cut}^{2}}{3\omega_{c}^{2}}\right)$
$\times\sum_{n,s}\left({\rm max}(n,n-s)\right)^{2}N^{\rm el}_{T}(n^{s})\left(1-N^{\rm el}_{T}(n^{s})\right).$ (51)
@@ Appendix E Conditions for neglecting the nonlinearity
$\Sigma_{n+1}^{\pm}$ $=\pm\frac{\hbar\omega_{r}}{2}\left(\sqrt{1+\left|\frac{2\lambda_{n+1,n}(\bm{0})}{\hbar\omega_{r}}\right|^{2}N_{0}}-1\right).$ (56)
@@
$n_{0}^{\rm crit}(n)\sim\frac{\varepsilon\hbar u}{2\pi ne^{2}\psi_{c}^{2}l_{H}^{3}}\sim\frac{m_{e}^{*}u}{\alpha\hbar l_{H}}.$ (59)
@@
$W{\simeq}\frac{cS_{\rm res}}{\sqrt{\varepsilon}}\int_{\hbar\omega_{\rm res}^{\rm min}(q)}^{\hbar\omega_{\rm res}^{\rm max}(q)}\!\frac{d^{3}q}{(2\pi\hbar)^{3}}N_{BE}\left({T_{\rm BB}}\right)\simeq$
$\simeq\frac{\varepsilon\omega_{\rm res}^{2}\Delta\omega_{\rm res}N_{\rm BE}(T_{\rm BB})}{2\pi^{2}c^{2}}<\varkappa\frac{\omega_{H}\Delta\omega}{c^{2}}Q$ (60)
@@
$n_{0}^{\rm BB}<\frac{W}{\varkappa_{0}+\varkappa}<Q\frac{\omega_{H}\delta\omega}{c^{2}}\ll\frac{m_{e}^{*}u}{\alpha\hbar l_{H}}\sim n_{0}^{\rm crit}.$ (61)
We can compare the critical density and photon density achieved by blackbody pumping: $n_{0}^{\rm BB}\ll n_{0}^{\rm crit}$. Thus, nonlinearities are safely negligible for our system.